Soil samples were collected at two areas surrounding an abandoned mine and analyzed for lead. At each area several samples were taken. The soil was extracted with acid, and the extract was analyzed using flame atomic absorption spectrometry. In Area A, Pb concentrations were 1.2, 1.0, 0.9, 1.4 mg/kg. In Area B, Pb concentrations were 0.7, 1.0, 0.5, 0.6, 0.4 mg/kg. Are these two areas significantly different from each other with Pb concentrations at 90% confidence level?


1
Expert's answer
2021-09-07T23:28:13-0400

Mean for Area A "=\frac{1.2+1.0+0.9+1.4}{4}=1.125mg/kg"


Mean for area B "=\frac{0.7+1.0+0.5+0.6+0.4}{5}=0.64"